### v1

Abstract
Streams of data have become the ubiquitous communication model on today’s Internet. For strong anonymous communication, this conflicts with the traditional notion of single, independent messages, as assumed e.g. by many mixnet designs. In this work, we investigate the anonymity factors that are inherent to stream communication. We introduce progressive pruning, a methodology suitable for estimating the anonymity level of streams. By mimicking an intersection attack, it captures the susceptibility of streams against traffic analysis attacks. We apply it to simulations of tailored examples of stream communication as well as to large-scale simulations of Tor using our novel TorFS simulator, finding that the stream length, the number of users, and how streams are distributed over the network have interdependent impacts on anonymity. Our work draws attention to challenges that need to be solved in order to provide strong anonymity for stream-based communication in the future.
Keywords:
anonymous communications, stream communication, anonymity metrics, intersection attacks, Tor, mixnets
1 Introduction
Anonymous communication is becoming increasingly important in today’s society, aiming to break sender-receiver linkability between communication partners. Like other types of online communication, it can either be carried out in the form of independent messages, or as streams, i.e. sequences of consecutive messages exchanged between the same sender-receiver pair. Today, data streams constitute the prevalent communication model underlying most practical applications on the Internet. In contrast, the traditional approach of anonymity systems regards independent messages as the base unit of communication. Attempts to break the users’ anonymity are typically carried out by a global or partially global adversaries as traffic analysis attacks. In such attacks, the adversary monitors the (encrypted) traffic and, based on the observed timing information, tries to correlate and thus link senders and receivers.
In this paper, we establish a more thorough understanding of how stream-based communication influences susceptibility to traffic analysis attacks. For this, we introduce progressive pruning, a methodology for estimating the anonymity level against intersection attacks, tailored to analyzing streams of communication. It works by combining the individual anonymity sets of messages as they occur over the progress of a stream, into a common pruned anonymity set for the whole stream. We argue that the perspective of progressive pruning is important in order to design future anonymity systems such that they can provide strong anonymity for stream communication. Progressive pruning enables estimation of the success chances of an intersection attack on the stream monitored by an attacker. For example, it can be used to compare the susceptibility to intersection attacks between different system variants.
We make use of this approach to derive general, asymptotic anonymity trends from stream properties. Our results indicate that anonymity is primarily influenced by the relative stream length among all streams, as well as the number of streams. We show that anonymity is distributed very unevenly and long streams implicitly take the role of cover traffic for others while experiencing poor anonymity themselves.
Our contributions in this work are the following:
  • We develop a methodology called progressive pruning (Section 3) to systematically evaluate the anonymity properties of streams, i.e. sequences of messages, as observed by an adversary who monitors traffic entering and leaving an ACN. We also discuss anonymity properties that can be derived.
  • We apply our optimized implementation of progressive pruning (ppcalc) to evaluate the anonymity impact of different stream properties based on synthetic network traces (Section 5).
  • We demonstrate how progressive pruning can be applied to a large-scale practical system by analyzing anonymity implications of the network topology in Tor. We carry out large-scale simulations of traffic flows in today’s Tor network as well as scaled variants (Section 6). For this, we introduce the Tor Flow Simulator (TorFS), a scalable alternative to the well-known TorPS [johnson2013usersrouted] that adds the ability to simulate traces of traffic flows.
2 Preliminaries
2.1 Background and Problem Statement
Figure 1: Anonymity of stream-based communication. For some user (yellow), an attacker would like to determine the communication partner (red or blue). A typical approach for de-anonymization would be an intersection attack.
Anonymous communication has been an active field of research for many years, starting from the original Chaumian mix design [chaum1981untraceable] and branching into various directions that aim to establish anonymity in online communication by following different strategies [edman2009anon]. One aspect that is of particular significance in this regard, but has received little attention, is the communication model employed by the different approaches. In general, these can be categorized into two distinct classes: On the one hand, the traditional approach assumes the notion of single independent messages that are transferred. This aligns well with the long time prevalent asynchronous communication based e.g. on email. On the other hand, anonymous communication can be regarded as streams of communication between senders and receivers as depicted in Figure 1. This understanding allows for synchronous, interactive communication and aligns more closely with modern online communication habits such as web browsing. These two different strategies can also be seen when looking at specific anonymous communication networks (ACNs). While mixnets such as Loopix [piotrowska2017loopix] generally transfer single messages over independent paths through the network, onion routing systems like Tor [dingledine2004tor] carry streams of messages (i.e. a stream of bytes, divided into equally-sized cells in the case of Tor). The two communication models—independent messages vs. stream-based communication—constitute fundamentally incompatible, orthogonal principles. This becomes apparent both when looking at the (performance) costs per message and especially when considering the security and anonymity implications. Mixnets like Loopix rely heavily on the fact that multiple messages between the same sender-receiver pair take independent paths through the network [piotrowska2017loopix].
Yet, we observe that there is a lack of research investigating the inherent factors that influence anonymity in stream-based anonymous communication. Many publications (implicitly) rely on the independence of transferred messages to derive anonymity properties—in particular, the unlinkability of multiple messages between the same sender-receiver pair. At the same time, when analyzing specific anonymity systems for interactive communication, shortcomings in the achieved anonymity are easily attributed to the low-latency nature of the system and lack of mixing, resulting in in the susceptibility against traffic analysis attacks. In contrast, we argue that it is crucial to understand the inherent anonymity implications of stream-based communication itself, independently of other system properties such as latency or cover traffic. This is because stream-based communication is at the core of modern online communication. Our work therefore aims to shed light on basic, inherent anonymity properties of stream-based communication that are relevant not only for onion routing systems but also to other ACNs aiming to support stream-based communication. One step towards understanding such relations is to measure or estimate the achieved anonymity. While a plethora of anonymity metrics exists for many different applications [DBLP:journals/csur/WagnerE18], these are not primarily useful for understanding the (asymptotic) implications of stream-based communication.
Our goal therefore is to quantify the anonymity level of streams in a way that helps understand inherent factors which increase or reduce anonymity.
2.2 Adversary Model
Discussing and quantifying anonymity only makes sense with regard to a concrete adversary model. In order to evaluate the anonymity properties of stream-based communication, we assume that an attacker can monitor both ends of the streams. That is, the attacker can observe when packets of each stream are sent and received. This setting therefore focuses on traffic analysis attacks in which the attacker compares temporal traffic patterns at both ends of the streams, trying to identify pairs of communication. In the context of stream-based communication the primary attack is an intersection attack [berthold2001disadvantages], which will therefore be the main concern of this work. The idea behind such attacks is to exploit the knowledge of multiple messages belonging together by intersecting their anonymity sets in order to identify the communication partners. Our adversary model is inspired by a global passive adversary, which appears to be a sensible assumption today [nsa-buys-data], and deliberately abstracts from any concrete underlying communication system to capture the anonymity properties inherent to streams.
2.3 Related Work
The long research history of anonymous communication begins in 1981, when Chaum introduced the concept of a mix [chaum1981untraceable]. Even in this early work, the central idea to achieve unlinkability revolves around the notion of independent messages that are uniformly sized and thus indistinguishable for an outside observer. This assumption has since become a core principle of mixnet designs, up to modern mixnet systems like Loopix [piotrowska2017loopix] and Nym [diaz2021nym].
The idea to anonymize streams of data was first followed in the context of ISDN telephony [pfitzmann1991isdn]. The system primarily exploited the fixed-rate nature of the ISDN network to achieve traffic analysis resistance. With the introduction of onion routing [syverson1997anonymous], the stream notion gained further attention, resulting in its adoption as a primary design principle in Tor [dingledine04tor]. Other low-latency anonymity networks have since continued to utilize this approach [chen2015hornet, DBLP:journals/popets/ChenP17].
However, streams of data being transferred in a low-latency fashion bring with them increased susceptibility to traffic analysis attacks, in particular intersection attacks [berthold2001disadvantages]. This is also the fundamental approach followed by our progressive pruning methodology, whereas our work aims specifically at making the impacts of specific stream properties comprehensible. A version of this attack that instead operates on the basis of probabilities has been put forward as the statistical disclosure attack. In general, the field of traffic analysis attacks has been the subject of many research studies. Correlation techniques that have been proposed range from simple approaches like packet counting [DBLP:conf/ih/BackMS01] through information-theoretic estimates like cross-correlation coefficients [levine2004timing] and mutual information [DBLP:conf/pet/ZhuFGBZ04] to advanced matching techniques based on deep learning [DBLP:conf/ccs/NasrBH18]. While these methods constitute relevant real-world attacks, they are not necessarily useful to understand the essence of anonymity within stream-based communication, as influenced e.g. by the length of streams or the number of simultaneous streams. We therefore propose an alternate methodology that supports establishing an understanding of how different inherent stream properties influence anonymity.
[…]
3 Progressive Pruning
3.1 Overview
We now introduce progressive pruning, a new methodology to analyze the anonymity implications of streams of data, as they are transferred via an ACN. We understand a stream as a reliable, unidirectional sequence of messages that are exchanged between the same pair of sender and receiver, over the same communication channel (e.g. through an ACN). Streams do not necessarily need to be order-preserving, however, messages need to be delivered within some reasonably bounded time. This transfer of streams of data comes with an increased vulnerability against intersection attacks. In such attacks, an attacker exploits the fact that multiple messages belong to the same communication partners by intersecting the anonymity set to reduce the number of candidates, keeping only candidates that are in line with all observations. Since the vulnerability to intersection attacks is what fundamentally differentiates streams from single messages, we design our method similarly. This way, we can identify anonymity effects that primarily stem from the stream characteristics, compared to the exchange of a single message.
The high-level goal of progressive pruning is as follows: Given a network trace of stream messages along with observation timestamps at its source and destination, we want to calculate an anonymity estimate for each of the contained streams. For this, progressive pruning builds upon the well-known notion of anonymity sets. However, it adds a time component: Initially, there may be a broad set of candidates, so the first message may have a large anonymity set. However, as communication proceeds over more messages, the anonymity set shrinks as candidates drop out. This may be the case because at some point their observed behavior may not be compatible with other observations. The fact that the set of candidates is pruned progressively as the message streams advance is what gives our methodology its name.
It should be noted that, as a consequence of the design, our anonymity metric does not provide an absolute value of system security or anonymity, which would be very hard or likely impossible. However, it can be used to analyze the relative changes to anonymity induced by different system parameters.
3.2 Detailed Description
Stream Model:
We assume that the input stream traces stem from a communication system of the following abstract notion: Unidirectional streams of messages are exchanged between senders and receivers. Messages can only be linked based on timing information and share a uniform size. We further assume that there is no message loss, and the delay of each message (the time between being sent and received) is bounded between $[D_{min},D_{max}]$. Depending on whether we want to evaluate the anonymity level of senders or receivers, either (1) senders send all their messages to the same receiver, or (2) receivers receive all their messages from the same sender. The attacker’s goal is to link senders and receivers. For clarity, we stick to the sender perspective from now on. Consequently, the attacker wants to find the corresponding receiver of each sender. The opposite direction is fully analogous. As our deliberately simple attacker represents a concrete attack instance, the resulting anonymity values can be interpreted as upper bounds on the achieved anonymity. Our assumptions regarding the stream model are not meant to mimic any existing real-world system, but rather allow investigating the inherent properties of stream communication.
Figure 2: Anonymity set calculation for individual messages of a sending stream. For each message from the sender (yellow), the set of potentially corresponding receiver messages is determined based on the lower and upper network delay bounds $D_{min}$ and $D_{max}$. Receivers of messages within this window (red, blue) may be communication partners of the sender. The anonymity set of the whole stream is then calculated by intersecting the individual message anonymity sets.
Terminology:
Let $S$ be the set of senders and $R$ the set of receivers. Accordingly, let $M_{S}$ be the set of messages sent by all senders, and $M_{R}$ the set of messages received by all receivers. $M_{S}[s]$ is the set of all messages sent by sender $s\in S$ and $M_{R}[r]$ is the set of all messages received by receiver $r\in R$. By $S[m]$, we denote the sender $s\in S$ of message $m$. Analogously, $R[m]$ is the receiver $r\in R$ of message $m$. The relationship between sent and received messages can be characterized as follows: $\forall m_{S}\in M_{S}:\exists m_{R}\in M_{R}:m_{S}=m_{R}$. It is the adversary’s goal to find this mapping, i.e. determine the receiver of a sent message, breaking sender-receiver unlinkabilty. Furthermore, let $T_{S}[m]$ be the time when message $m$ was sent, and $T_{R}[m]$ the time when message $m$ was received.
Anonymity set of a single message:
We first define the anonymity set for a single message before extending it to multiple messages afterwards (cf. Figure 2). To do so, we introduce the notion of a message-message anonymity set $\Psi_{MM}$. For each sent message $m_{S}$, this is the set of received messages $m_{R}$ that $m_{S}$ could correspond to, because they were received within the timeframe assumed for message delays. We therefore define it as follows:
$\Psi_{MM}[m_{S}]$ =$=\{m_{R}$ $\in M_{R}\mid T_{R}[m_{R}]\in[T_{S}[m_{S}]+D_{min}~{},~{}T_{S}[m_{S}]+D_{max}]\}$
Based on this, it is straightforward to define the message-receiver anonymity set $\Psi_{MR}$, of a message, denoting the set of potential receiver candidates, a subset of $R$:
$\Psi_{MR}[m_{S}]=\{R[m_{R}]~{}\forall~{}m_{R}\in\Psi_{MM}[m_{S}]\}$
Anonymity set of multiple messages:
We now extend these anonymity set definitions to multiple messages from the same sender (a stream). By doing so, we want to capture the anonymity set of senders as a whole, taking all their messages into account. The simplest approach would be to simply intersect the message-receiver anonymity sets for all messages from a sender: $\bigcap_{m_{S}\in M_{S}[s]}\Psi_{MR}[m_{S}]$. However, this falls short of the fact that the adversary can trivially also exploit the number of messages to differentiate between candidates. For example, if a sender sent 10 messages in a short period of time, but a receiver received only one message that falls into the message-message anonymity sets of all sent messages, the simple intersection approach would not exclude that receiver although it obviously cannot be a candidate because it received too few messages.
We solve this issue by formulating the following assignment problem: Let $s$ be a sender of messages $M_{S}[s]$ and $r$ a receiver of messages $M_{R}[r]$. Then, $r$ can only be a candidate as the receiver of the messages from $s$, if there exists a mapping $\sigma:M_{S}[s]\to M_{R}[r]$ that assigns each sent message to a corresponding received one and meets the following two requirements: (1) If $\sigma(m_{S})=m_{R}$, then $m_{R}\in\Psi_{MM}[m_{S}]$ (the messages potentially correspond with respect to the delay constraints), and (2) $\sigma$ is injective (each received message is assigned at most once).
Given the sets of sent and received messages, it is possible to determine whether such a mapping exists, e.g. in a brute-force manner. Moreover, in Section 3.4 we will present an efficient algorithm to solve this task in linear time. If a mapping that meets the the aforementioned requirements can be found, we say that $r$ is a candidate for $s$. Consequently, we can define the sender-receiver anonymity set of sender $s$ as follows:
$\Psi_{SR}[s]=\{r\in R\mid\text{$r$ is a candidate for $s$}\}$
This now denotes the final anonymity set usable to estimate the anonymity set of a sender. The anonymity set size $A(s)=|\Psi_{SR}[s]|$ can consequently be regarded as a measure of the achieved anonymity. If $A(s)=1$, then $s$ can be de-anonymized by an adversary that observes sent and received messages. If $A(s)>1$, then the simple window-based attack we modeled would not be successful because multiple candidates remain.
3.3 Model Discussion
The utility of progressive pruning is based on the fact that it couples the anonymity sets of a sender’s individual messages to achieve an estimate of the sender’s anonymity. It is therefore suitable for evaluating the anonymity of stream-based communication, as opposed to single messages.
Intuitively, we can understand the presented construction as the anonymity set of a sender $s$ over time. When taking into account only a single message, then all receivers that have received a message in the respective timeframe are possible communication partners of $s$. Additional messages in the stream, however, add additional constraints that can be exploited by the adversary. As such, more messages will likely result in a shrinking anonymity set (if the streams of communication are not synchronized). At the same time, the number of observed messages that are necessary to break anonymity can be regarded as an indicator for the costs an adversary has to invest for the attack.
Another observation that can be made is that the number of messages sent by a sender is highly relevant for the resulting anonymity set. In particular, if the streams differ in their number of messages, then the stream length is useful for an adversary to differentiate the senders. While this comes at no surprise, it has an interesting implication: A sender that sends more data than anyone else does not retain any anonymity at all. In general, the shorter the communication is in comparison to the others, the more likely it is that a sender can “hide” among the other senders. We formally verified this important relationship and present the analysis in the appendix. Additionally, the absolute latency value of messages does not directly make a difference. Instead, an attacker only benefits from a low deviation of latency (jitter) because it makes the traffic more predictable, allowing use of a smaller window.
3.4 Algorithm: Matching Candidate Messages
Here we present an efficient algorithm that can be used to determine whether a receiver $r\in R$ is a candidate for sender $s\in S$, based on their sets of sent and received messages, as defined in Section 3.2. Put differently, the algorithm decides whether a suitable assignment $\sigma$ exists, mapping the messages sent by $s$ to valid messages received by $r$.
We exploit the fact that the exact mapping is not needed, but it is sufficient to determine if one exists. For each sent message $m_{S}\in M_{S}[s]$, we first determine the potential candidate messages in $r$:
$\psi(m_{S},r)=\Psi_{MM}[m_{S}]~{}\cap~{}M_{R}[r]$
Without loss of generality, we then iterate over the sent messages $m_{S}\in M_{S}[s]$, ordered by their receive timestamp. If the messages from a sender are always delivered in order, this corresponds to the observed send sequence. In the special case that we evaluate a complete “white-box” network trace containing the ground truth message mapping, we can also determine that order. In practice, if messages are allowed to be re-ordered, an attacker would not know the order and might have to choose a less efficient algorithm. This, however, is solely a matter of computational efficiency for an attacker.
During iteration, we need to decide whether $m_{S}$ can be mapped to a received message. We do so by comparing $\psi(m_{S},r)$ to the (previous) message anonymity set $\psi(m_{S}^{\prime},r)$ of the previous message $m_{S}^{\prime}$. $m_{S}$ can mapped if and only if there is at least one valid candidate message. The number of candidates (cand) is determined as the sum of leftover unused candidate messages from the previous anonymity set $\psi(m_{S}^{\prime},r)$ that are also present in $\psi(m_{S},r)$, which we call overlap, plus the number of new messages $|\psi(m_{S},r)\setminus\psi(m_{S}^{\prime},r)|$, which we call new. If cand > 0, then $m_{S}$ can be mapped and the next message continues with cand - 1. If for all messages, the number of candidates is non-zero, then a suitable assignment $\sigma$ exists. Otherwise, this is not the case.
3.5 Implementation: ppcalc
We implement progressive pruning and make it available as an open-source project called ppcalc  [^1] https://github.com/cdoepmann/ppcalc, implemented in the Rust language. ppcalc consumes network traces consisting of messages between senders and receivers, and their timestamps. It outputs the anonymity set of each stream. It processes plain CSV data and can thus easily be reused for other research questions. It supports the calculation of sender anonymity as well as receiver anonymity. During the implementation of ppcalc, great care was taken to achieve high scalability that enables ppcalc to process communication traces not only from relatively small, synthetic scenarios as considered in this section, but also traces from real-world networks like Tor, at scale. For the experiments we will later present in Section 6, ppcalc successfully managed to process network traces of approximately 720 GB within 74 hours on a 12 CPU cores machine.
4 Understanding Anonymity Over Time
Progressive pruning, as presented in Section 3.2, explicitly focuses on the evaluation of anonymity within stream-based communication. That is, it takes into account how the anonymity set changes over the course of a data transfer that consists of a sequence of messages. In order to understand the implications of this measurement approach, we therefore analyze how anonymity changes with an ongoing stream of messages. This can also be understood as a development over time as the single messages are generally sent sequentially. Note that progressive pruning does not require any specific temporal characteristics but works the same if all messages in a stream are sent simultaneously.
For this first analysis, we rely purely on stochastical methods in order to learn about the time behavior that is inherent to our model. In order to do so, we first consider the following very simple model of communication over an anonymous communication network: For now, we assume that all users send their messages at the same rate. This assumption allows us to discretize the system into synchronized rounds, assuming that one round corresponds to the transmission of exactly one message by each user that is active during that round. Although this is a simplification, we argue that is still close to the overall model behavior due to its window semantics. Accordingly, the window size that defines which messages belong to the anonymity set corresponds to one round. Making use of the terminology from Section 3.2, messages that are sent in the same round form a common message-message anonymity set. More intuitively, senders belong to the same anonymity set if they send data at the same time (within the same window/round). This matches our intuitive understanding of traffic analysis, as users sending messages at clearly different times are easy to tell apart. For simplicity, we further assume that all senders transfer the same number $L$ of messages, in $L$ consecutive rounds.
Let us consider a single sender $s$ within a busy anonymity system, starting to send its messages at round 00. Then, the transfer will last $L$ rounds. Moreover, let $N$ be the overall number of (independent) senders active in the system at this first round. For each round $t$, we want to find the anonymity set size $|\Psi_{SR}[s]|$, which we abbreviate as $A_{t}$ here. Obviously, $A_{0}=N$. Since our sender $s$ starts sending while the system is already in use by other senders, those senders will finish sending their messages at different times. Let $E_{i}$ be the time when sender $i$’s transfer ends (more precisely, the first round in which sender $i$ does not send a message anymore). Then, $E_{s}=L$ because $s$ sends for $L$ rounds and starts at round 00.
In this scenario, the anonymity set will only shrink when other senders stop sending at some time. Note that users starting to send later do not have an impact on the anonymity set because they have not been candidates before and thus cannot be valid anonymity set members later. The anonymity set cannot grow over time because it corresponds to the intersection of observations over time. The anonymity set size is therefore determined by the number of senders that have stopped sending. Let $\bar{A}_{t}$ be the number of senders that have already finished their transfer before round $t$. Then the anonymity set size is $A_{t}=N-\bar{A}_{t}$, and $\bar{A}_{0}=0$.
Without loss of generality, let us assume that $E_{s}$ corresponds to $E_{0}$, simplifying the following formulations. Since all senders transfer the same number $L$ of messages, and are independent from each other, we can calculate the probabilities for certain end times as follows:
$P(E_{i}=1)=\dots=P(E_{i}=L)=\frac{1}{L}~{}~{}~{}\text{for $0<i\leq N$}$
Consequently, we can accumulate the probabilities as follows (note that the events $\{E_{i}=t\}$ are all disjoint):
$P(E_{i}\leq 0)=0\qquad P(E_{i}\leq L)=1\qquad P(E_{i}\leq t)=t\cdot\frac{1}{L}$
The relationship between $\bar{A}_{t}$ and the different $E_{i}$ is as follows: For a fixed round $t$, $\bar{A}_{t}=t$ iff. $E_{i}\leq t$ for $x$ different $i$, and $E_{i}>t$ for the remaining $(N-x)$ $E_{i}$. Due to the fact that the events $\{E_{i}\leq t\}$ are all independent of each other, we can model each of them as a Bernoulli experiment with success probability $p=t\cdot\frac{1}{L}$. For each round $t$, $\bar{A}_{t}$ is therefore binomially distributed with $n=N$ and $p=t\cdot\frac{1}{L}$:
$P(\bar{A}_{t}=x)=\binom{n}{x}\cdot p^{x}\cdot(1-p)^{n-x}$
Consequently, the anonymity set size is distributed as follows:
$P(A_{t}=x)=P(\bar{A}_{t}=N-x)=\binom{N}{N-x}\cdot p^{N-x}\cdot(1-p)^{x}\qquad\square$
We can draw interesting insights about the anonymity set size as modeled by our progressive pruning. In particular, if we calculate the expected value for the anonymity set size in round $t$, we get:
$E[A_{t}]=E[N-\bar{A}_{t}]=N-E[\bar{A}_{t}]=N-n\cdot p=N-N\cdot\left(t\cdot\frac{1}{L}\right)$
As a consequence, we can see that the expected anonymity set size decreases linearly/anti-proportionally over time. Put differently, the longer the transfer lasts, the weaker the achieved anonymity gets. This observation clearly underlines the significance of the data stream length: The shorter the communication is, the more likely it is to hide among other senders. Also note that the simple system setup described here (same-length senders with independent start times) is equivalent in principle to another relevant scenario to consider in which all senders start sending at the same time but have independently selected transfer lengths.
5 Anonymity Factors for Streams
{subcaptionblock}
.3 Refer to caption {subcaptionblock}.3 Refer to caption {subcaptionblock}.3 Refer to caption
Figure 3:
Figure 4:
Figure 5:
Figure 6: Influence of different stream properties on stream anonymity set size.
We now want to gain initial insight on the general factors that influence anonymity in the case of stream communication. By doing so, we aim to verify common assumptions and expectations about the development of anonymity in stream communication. Note that this investigation complements existing work like [wright2003defending] and [johnson2013usersrouted] in that these focus on anonymity over time due to user behavior, while we consider anonymity properties of the streams themselves.
Our general approach is as follows: We simulate communication between abstract senders and receivers with different scenario parameters. This simulation of communication streams is carried out on a high abstraction level—the communication properties are defined only by general parameters like stream length, latency, and number of senders and receivers. We then apply progressive pruning and use the calculated anonymity set sizes as estimates for anonymity. By altering the scenario parameters, we can evaluate the relative influence of each parameter on anonymity. Without loss of generality, we focus on sender anonymity here. Therefore, each sender sends all of their messages to one receiver, and the anonymity set of a sender corresponds to the set of candidate receivers as observed by an attacker.
As our base scenario, we choose 2,000 senders, starting their transfer within a time span of 60 seconds. This number is loosely inspired by the typical 3 million daily users in Tor. Each sender transmits 2,314 KB of data, corresponding to the average size of a website. [^2] https://almanac.httparchive.org/en/2022/page-weight#request-bytes Senders transmit at a data rate of 10 Mbit/s, sending to a total of 400 receivers, chosen uniformly at random. The network latency is set to 210 ms, a typical value found in the real Tor network. [^3] https://metrics.torproject.org/onionperf-latencies.csv?start=2023-05-05&end=2023-08-03&server=public The analysis window that bounds network latency for anonymity calculation in progressive pruning is set to ±$\pm 100$ ms around this value. For each of the parameters, we carry out a series of simulations in which we only vary this parameter.
Our experiments reveal and verify the following relationships: Firstly, the length of a stream, in relation to other streams in the network, is of crucial importance for the achieved anonymity (cf. Figure 6). While short streams can achieve large anonymity sets, longer streams serve as cover traffic for the whole network but lose their own anonymity. This underlines the effectiveness of cover traffic, either implicitly through a large user base, or explicitly using designated cover traffic. Secondly, a low number of receivers compared to the number of senders is beneficial for anonymity because it helps streams to blend in with each other (Figure 6). And thirdly, homogeneity of streams is also beneficial for anonymity. (Figure 6).
While these results come as no surprise, they express the soundness and utility of predictive pruning to analyze anonymity properties of streams, also verifying these relationships based on a well-defined stream model.
6 Case Study: Comparing Tor Instances
In the previous section, we demonstrated the behavior of progressive pruning for the analysis of stream-based communication, and inferred general factors that are relevant for the achieved anonymity. The considered scenarios, however, were synthetic examples suitable primarily for investigating and verifying our metric as well as general, abstract trends. In contrast, we now go further and aim to apply our analysis approach to the most widely used anonymity network as of today—the Tor network. By doing so, we want to give a practical example of how progressive pruning can be used to research the anonymity properties of large-scale anonymity systems providing stream communication.
As laid out in Section 3, progressive pruning is not suitable for quantifying anonymity as an absolute value, but rather as a relative measure for comparing anonymity. We therefore refrain from quantifying “the anonymity level of Tor”. The quantification can, however, be used to compare different instances of the network, hinting at deviations in achieved anonymity. We use this approach to investigate the implications of one of the future changes Tor will face, namely network growth.
Over time, the Tor network has grown significantly and it is expected to continue this trend in the future [DBLP:conf/IEEEares/DopmannT23]. While this development indicates the increasing significance and the growing user base of Tor, it does not come without challenges. Upcoming issues should be spotted early, so they can be taken into consideration when designing future network iterations. In [DBLP:conf/IEEEares/DopmannT23], the authors introduced a methodology to investigate upscaled instances of the Tor network, with regard to different topological properties, also finding initial evidence that purely horizontal growth of the Tor network can lead to suboptimal performance characteristics. We now aim to leverage our progressive pruning methodology to analyze the scaling impact on anonymity provided by the network.
6.1 Methodology
Our general approach is to estimate the achieved anonymity based on simulations of the network. For this, we first simulate the Tor network in order to obtain a set of circuits that were created, along with traces of the exchanged messages (cells) and their timestamps. We then compute the anonymity level according to progressive pruning. This process is first carried out for the original (unscaled) Tor network. Using torsynth [DBLP:conf/IEEEares/DopmannT23], we then generate two synthetic but representative Tor networks that were scaled either vertically or horizontally by +100%. We also evaluate these networks using the aforementioned workflow and then compare the achieved anonymity values. In order to avoid sampling artifacts, we refrain from down-sampling the networks for simulation. Instead, we carry out all experiments at scale, which constitutes a nontrivial engineering challenge to scalability of the involved software components. We base our experiments on the network consensus from 2023-10-01 at 00:00, and simulate one hour of full network activity.
Tor Flow Simulation:
Given a specific snapshot of the Tor network (consensus), the network simulation step needs to generate a trace of network activity within the Tor network. In particular, this involves generating circuits and using them to transport data cells as would be the case in the real network.
There are different approaches towards simulating the Tor network, trading between computational effort and realism. The most realistic simulations can be carried out using the shadow simulator which is able to run the unmodified Tor client. However, running such simulations at scale 1:1 (or even 2:1 in the case of up-scaled networks) was infeasible. While shadow enables a very fine-grained simulation to investigate performance and other networking effects, we are particularly interested in how circuits and cells are distributed across the network because these are the factors for anonymity as considered by our metric. We therefore choose a more abstract simulation approach, similar to the Tor path simulator (TorPS [johnson2013usersrouted]). TorPS simulates circuit generation by Tor users over time and has been used before to evaluate how anonymity changes when users construct many consecutive circuits. However, TorPS suffers from severe scalability issues. Furthermore, it does not simulate user behavior at the packet level and thus cannot be used to generate the network traces we require.
We therefore propose a novel simulator software which we call Tor Flow Simulator (TorFS) and make available as an open source project [^4] https://github.com/cdoepmann/torfs, TorFS inherits much of TorPS’s functionality. However, it is implemented in the Rust programming language with a focus on scalability, and adds the missing pieces needed to simulate Tor user behavior at the packet level. Most importantly, it applies the modeled circuit selection rules as well as the users’ circuit management. It also supersedes TorPS’s simplistic user model, replacing it with the state-of-the-art model introduced in [DBLP:conf/ccs/JansenTH18] and used by other publications in the field since then [DBLP:journals/popets/PahlAKE23, DBLP:conf/ccs/RochetWJMP20, DBLP:conf/uss/JansenVS19]. As in [DBLP:conf/ccs/JansenTH18], user behavior is now modeled as a probabilistic process (an exponential distribution) with parameters derived from real-world measurements on the Tor network. Moreover, while TorPS simulates only the generation and management of circuits, TorFS also simulates the circuits themselves to generate network traces of the individual messages (cells). For this, it again makes use of the probabilistic model from [DBLP:conf/ccs/JansenTH18]. Consequently, TorFS uses the stream model as well as the packet model from [DBLP:conf/ccs/JansenTH18] which are Markov chains to mimic real traffic behavior as observed on the real Tor network. At the same time, TorFS does not constitute a full-fledged network simulator like shadow or ns-3. It abstracts from networking effects in that cells entering the network are output after a constant latency at their destination. This makes it feasible to simulate the network at scale. Moreover, we argue that this is in line with our approach of focusing on the inherent anonymity effects within the Tor network as influenced by its topology.
TorFS outputs message traces ready for analysis with ppcalc. We only focus on traffic that is directed from the Tor exit to the client because this constitutes the majority of Tor traffic. We apply the following mapping from simulated Tor behavior to the abstract message model from Section 3: Each cell that occurs in the simulation corresponds to a message in the abstract model. Multiple cells that form a data stream in the simulation also constitute a single stream of messages in the sense of our abstract model. The abstract sender of these messages is the according simulated exit. For each stream, we define an abstract receiver as the entity that receives all of this stream’s messages. This way, the stream-based anonymity can be evaluated independently for each stream in the network.
Anonymity Computation:
After generating message traces of the Tor network using TorFS, we analyze these using progressive pruning in order to estimate the anonymity level. For this task, we again used ppcalc (cf. Section 3.5). As laid out before, the anonymity estimate can be calculated in either one of two possible “directions”. In our current experiment, we have a fixed (relatively small) set of exit relays that are sending data to many clients. Due to this 1:n relationship, it makes sense to calculate anonymity from the receiver perspective. That is, for each receiver (client), we calculate the set of potential senders (exits) from which their data may originate. Analyzing anonymity in the reverse direction (the set of potential receivers per sender) is not useful in this scenario due to the fact that exits are sending to many clients at the same time.
6.2 Results and Discussion
Table 1: Anonymity results from simulations of an original Tor consensus, Tor scaled vertically by +100%, and Tor scaled horizontally by +100%.
Original Vertical Horizontal
Fully deanonymized streams 795/26,407,264 41/52,808,426 1096/52,773,747
Avg/median anonymity set size 1,241.4 / 1,440 1,610.8 / 1,784 2,481.9 / 2,882
Stddev of anonymity set size 594.4 416.5 1,186.0
{subcaptionblock}
.45 Refer to caption {subcaptionblock}.45 Refer to caption
Figure 7: Orig. (Refer to caption), vert. (Refer to caption), horiz. (Refer to caption)
Figure 8: Original
{subcaptionblock}
.45 Refer to caption {subcaptionblock}.45 Refer to caption
Figure 9: Vertically scaled
Figure 10: Horizontally scaled
Figure 11: (a) Overall anonymity distribution of streams per scenario, (b)-(d) Anonymity distribution of streams, grouped by stream length per scenario (from percentiles 0-10/black, to percentiles 90-100/orange).
Our analysis is based on the streams’ anonymity set sizes and their distribution across the network. Table 1 displays fundamental statistics about the anonymity in each simulated network. As a first observation, we can see that, on average, the anonymity set sizes are relatively large. In the case of the unmodified Tor network, e.g., streams still remain with an average of more than 1,200 sender candidates (out of a total number of approximately 2,400 exits). It is noteworthy that, with regard to the simple window-based attack progressive pruning mimics, anonymity of the vast majority of streams is not trivially broken. This observation is supported by the fact that—looking again at the original Tor network—out of 26 million streams, less than 1,000 are fully de-anonymized (resulting in an anonymity set size of 1).
In order to better understand how anonymity is distributed within the network, we consider CDF plots of the achieved anonymity set sizes, displayed in Figure 11. Looking at the data from the original Tor network first, it becomes immediately apparent that anonymity is distributed very unevenly between the streams. Note that, if all streams had the same anonymity set size, the CDF plots would exhibit a single vertical line. In contrast, the simulation results show that there are many streams with very little anonymity remaining, just like there are many streams with rather strong anonymity remaining and many streams in between. While a result like this was to be expected due to the fact that Tor has no means of ensuring a certain level of anonymity, it still shows a surprisingly clear weakness of the system because users cannot reasonably expect a certain level of anonymity, as far as the stream-based anonymity degradation over time is concerned.
Having seen in Section 5 that a stream’s anonymity is heavily influenced by its length, we next investigate the relationship between stream lengths and anonymity in the simulated networks. Figure 11 contains CDF plots of the achieved anonymity set sizes again, but subdivides the streams into quantiles based on their length (number of cells). The resulting anonymity distributions clearly point out that stream length is again a decisive factor for the resulting anonymity. Long-lasting streams are massively disadvantaged and generally achieve only little anonymity. In contrast, short streams are much harder to de-anonymize and have anonymity sets that are multiple times larger. These results again support our observations from Section 5: Long streams suffer from reduced anonymity but act as cover traffic providing stronger anonymity for shorter streams.
Anonymity in Scaled Networks
We leverage our simulation and analysis workflow to evaluate not only the Tor network as is today but also variations of the network that could result from future network growth of different characteristics. Our choice to evaluate scaled Tor networks is motivated by two goals: Firstly, it allows us to research the influence of network topology on the achieved anonymity. And secondly, understanding the trade-offs of different topological properties may provide important insights on how to shape and prepare for future network evolution. In addition to a snapshot of today’s Tor network, we therefore carried out simulations of the Tor network being scaled by +100% vertically as well as by +100% horizontally. By “vertical” growth we mean that the bandwidth capacity of each relay is increased accordingly. Likewise, “horizontal” growth is achieved by adding new, additional relays to the existing network. The scaling methodology of torsynth is detailed in [DBLP:conf/IEEEares/DopmannT23]. For simulations, we also grow the number of active users on the network by +100% so the networks differ only in their topology, and not in the amount of available resources per user.
We first consider the case of horizontal growth. Table 1 and Figure 11 show that increasing the number of relays has a strong impact on anonymity as defined by progressive pruning. In fact, with double the number of relays, the average anonymity size is also exactly doubled. This trend is intuitively reasonable because doubling the number of relays also implies doubling the number of available exits which are the potential anonymity set candidates. However, it is interesting to see that this increase of average anonymity set size appears to be exactly proportional to the number of exit relays. This is somewhat surprising because, essentially, the number of exits only directly defines the maximum anonymity set size which wouldn’t necessarily have to correspond linearly to the average anonymity set size. It is noteworthy that the standard deviation also doubles exactly.
Looking at the vertically upscaled network instead paints a rather different picture. Due to the fact that the number of exits—and thus anonymity set candidates—has not increased, the average anonymity set size has not doubled as in the case of horizontal growth. While this was to be expected, the vertical growth (and the increased load) of the network still results in a clear increase of anonymity by approximately 30%. Combined with our previous observations, this makes clear that there is no single parameter that determines the anonymity level. Instead, we see that there are two primary factors for anonymity as defined by our metric: topology on the one hand (determining the maximum number of potential anonymity set members) and traffic load on the other. Also note that, compared to the original Tor network, the vertically scaled instance exhibits less variation in anonymity set size among the streams which is in contrast to the previously observed increase in the case of horizontal growth, and which might be a desirable system property.
All in all, we can conclude from the simulations of the Tor network which we carried out that there are two fundamental factors for anonymity in the network: Firstly, an increased number of relays also increases the number of (topological) candidates for the anonymity set. And secondly, apart from the topological factor, the mere increase in network load also improves anonymity, albeit much less effectively.
7 Conclusion
In this work, we analyzed the inherent susceptibility of stream-based anonymous communication to traffic analysis attacks. In particular, we focused on intersection attacks because these constitute the class of attacks which streams are fundamentally vulnerable against, compared to independent messages. We introduced progressive pruning, a methodology to quantify the anonymity level of streams, with regard to the effectiveness of an intersection attack. With that, we contribute a new perspective: Instead of regarding the susceptibility to intersection attacks as a binary property, we enable a more differentiated view. The ability to compare the anonymity achieved by different system variants can be used for designing new ACNs, taking attackers’ cost into account.
We carried out simulations to investigate the impact of different stream properties in tailored synthetic scenarios. Moreover, we demonstrated how progressive pruning can be applied to large systems like the Tor network, estimating anonymity factors with regard to network topology. Our results indicate that the susceptibility towards traffic analysis attacks of streams is primarily dependent on the length of streams, relative to each other. Longer streams suffer from worse anonymity because they can more easily be distinguished from others by considering their number of messages. At the same time, they serve as cover traffic, increasing anonymity for shorter streams. When looking at Tor as an example of stream-based anonymity networks, we observe different influences when scaling the network. On the one hand, scaling the network horizontally (adding relays) helps increase anonymity set sizes in the context of our anonymity estimate that takes the perspective of a global adversary who monitors incoming and outgoing traffic of the anonymity system. On the other hand, when considering a vertically scaled network, we can see that an increased number of users is beneficial for anonymity.
We argue that stream-based communication is of high practical relevance for online privacy and should therefore be a goal when designing future anonymity systems. While our work shows that it has inherent, negative consequences for anonymity due to being vulnerable to intersection attacks, our results may serve as a starting point in working around these limitations with the vision of combining stream-based communication with strong anonymity.
[…]

### v2

@@
 Abstract
-Streams of data have become the ubiquitous communication model on today’s Internet.
-For strong anonymous communication, this conflicts with the traditional notion of single, independent messages, as assumed e.g.
-by many mixnet designs.
-In this work, we investigate the anonymity factors that are inherent to stream communication.
-We introduce progressive pruning, a methodology suitable for estimating the anonymity level of streams.
-By mimicking an intersection attack, it captures the susceptibility of streams against traffic analysis attacks.
-We apply it to simulations of tailored examples of stream communication as well as to large-scale simulations of Tor using our novel TorFS simulator, finding that the stream length, the number of users, and how streams are distributed over the network have interdependent impacts on anonymity.
-Our work draws attention to challenges that need to be solved in order to provide strong anonymity for stream-based communication in the future.
+Stream-based communication dominates today’s Internet, posing unique challenges for anonymous communication networks (ACNs).
+Traditionally designed for independent messages, ACNs struggle to account for the inherent vulnerabilities of streams, such as susceptibility to intersection attacks.
+In this work, we address this gap and introduce progressive pruning, a novel methodology for quantifying the susceptibility to intersection attacks.
+Progressive pruning quantifies and monitors anonymity sets over time, providing an assessment of an adversary’s success in correlating senders and receivers.
+We leverage this methodology to analyze synthetic scenarios and large-scale simulations of the Tor network using our newly developed TorFS simulator.
+Our findings reveal that anonymity is significantly influenced by stream length, user population, and stream distribution across the network.
+These insights highlight critical design challenges for future ACNs seeking to safeguard stream-based communication against traffic analysis attacks.
 Keywords:
-anonymous communications, stream communication, anonymity metrics, intersection attacks, Tor, mixnets
+anonymous communications, intersection attacks, Tor
 1 Introduction
-Anonymous communication is becoming increasingly important in today’s society, aiming to break sender-receiver linkability between communication partners.
-Like other types of online communication, it can either be carried out in the form of independent messages, or as streams, i.e.
-sequences of consecutive messages exchanged between the same sender-receiver pair.
-Today, data streams constitute the prevalent communication model underlying most practical applications on the Internet.
-In contrast, the traditional approach of anonymity systems regards independent messages as the base unit of communication.
-Attempts to break the users’ anonymity are typically carried out by a global or partially global adversaries as traffic analysis attacks.
-In such attacks, the adversary monitors the (encrypted) traffic and, based on the observed timing information, tries to correlate and thus link senders and receivers.
-In this paper, we establish a more thorough understanding of how stream-based communication influences susceptibility to traffic analysis attacks.
-For this, we introduce progressive pruning, a methodology for estimating the anonymity level against intersection attacks, tailored to analyzing streams of communication.
-It works by combining the individual anonymity sets of messages as they occur over the progress of a stream, into a common pruned anonymity set for the whole stream.
-We argue that the perspective of progressive pruning is important in order to design future anonymity systems such that they can provide strong anonymity for stream communication.
-Progressive pruning enables estimation of the success chances of an intersection attack on the stream monitored by an attacker.
-For example, it can be used to compare the susceptibility to intersection attacks between different system variants.
-We make use of this approach to derive general, asymptotic anonymity trends from stream properties.
+Anonymous communication is becoming increasingly vital in today’s society, as it aims to prevent the correlation of senders and receivers in communication networks.
+While early anonymous communication networks (ACNs) focused on high-latency networks operating with independent messages, the dominant model today involves low-latency networks designed to transport streams of data.
+A notable example is the Tor network [dingledine04tor].
+However, a fundamental vulnerability of such systems lies in traffic analysis attacks.
+In these attacks, adversaries monitor (encrypted) traffic and use observed timing information to correlate and link senders with receivers.
+Recent incidents, such as the German Federal Criminal Police Office (BKA) successfully executing a timing attack [panorama-boystown], highlight the pressing need to address these vulnerabilities.
+Dismissing such attacks as out of scope is no longer a viable option.
+In this paper, we establish a deeper understanding of how stream-based communication impacts susceptibility to traffic analysis.
+For this, we introduce progressive pruning, a methodology for quantifying anonymity sets under intersection attacks, specifically designed for analyzing streams of communication.
+Progressive pruning works by combining anonymity sets of individual messages as they occur over the course of a stream, into a common pruned anonymity set for the entire stream.
+This approach enables estimating an attacker’s success chances when performing an intersection attack on a monitored stream.
+For instance, progressive pruning can be used to compare the susceptibility of different system variants to such attacks.
+We argue that adopting this perspective is crucial for designing future anonymity systems capable of providing robust anonymity for stream-based communication.
+We make use of progressive pruning to derive general, asymptotic anonymity trends from stream properties.
 Our results indicate that anonymity is primarily influenced by the relative stream length among all streams, as well as the number of streams.
-We show that anonymity is distributed very unevenly and long streams implicitly take the role of cover traffic for others while experiencing poor anonymity themselves.
+We show that long streams implicitly take the role of cover traffic for others while experiencing poor anonymity themselves.
+Applying progressive pruning to Tor as a real-world network, we find that anonymity is distributed very unevenly.
 Our contributions in this work are the following:
-  • We develop a methodology called progressive pruning (Section 3) to systematically evaluate the anonymity properties of streams, i.e.
-sequences of messages, as observed by an adversary who monitors traffic entering and leaving an ACN.
+  • We develop a methodology called progressive pruning (Section 3) to systematically evaluate the anonymity properties of streams, i.e., sequences of messages, as observed by an adversary who monitors traffic entering and leaving an ACN.
 We also discuss anonymity properties that can be derived.
@@
 For this, we introduce the Tor Flow Simulator (TorFS), a scalable alternative to the well-known TorPS [johnson2013usersrouted] that adds the ability to simulate traces of traffic flows.
-2 Preliminaries
-2.1 Background and Problem Statement
-Figure 1: Anonymity of stream-based communication.
-For some user (yellow), an attacker would like to determine the communication partner (red or blue).
-A typical approach for de-anonymization would be an intersection attack.
-Anonymous communication has been an active field of research for many years, starting from the original Chaumian mix design [chaum1981untraceable] and branching into various directions that aim to establish anonymity in online communication by following different strategies [edman2009anon].
-One aspect that is of particular significance in this regard, but has received little attention, is the communication model employed by the different approaches.
-In general, these can be categorized into two distinct classes: On the one hand, the traditional approach assumes the notion of single independent messages that are transferred.
-This aligns well with the long time prevalent asynchronous communication based e.g.
-on email.
-On the other hand, anonymous communication can be regarded as streams of communication between senders and receivers as depicted in Figure 1.
-This understanding allows for synchronous, interactive communication and aligns more closely with modern online communication habits such as web browsing.
-These two different strategies can also be seen when looking at specific anonymous communication networks (ACNs).
-While mixnets such as Loopix [piotrowska2017loopix] generally transfer single messages over independent paths through the network, onion routing systems like Tor [dingledine2004tor] carry streams of messages (i.e.
-a stream of bytes, divided into equally-sized cells in the case of Tor).
-The two communication models—independent messages vs.
-stream-based communication—constitute fundamentally incompatible, orthogonal principles.
-This becomes apparent both when looking at the (performance) costs per message and especially when considering the security and anonymity implications.
-Mixnets like Loopix rely heavily on the fact that multiple messages between the same sender-receiver pair take independent paths through the network [piotrowska2017loopix].
-Yet, we observe that there is a lack of research investigating the inherent factors that influence anonymity in stream-based anonymous communication.
-Many publications (implicitly) rely on the independence of transferred messages to derive anonymity properties—in particular, the unlinkability of multiple messages between the same sender-receiver pair.
-At the same time, when analyzing specific anonymity systems for interactive communication, shortcomings in the achieved anonymity are easily attributed to the low-latency nature of the system and lack of mixing, resulting in in the susceptibility against traffic analysis attacks.
-In contrast, we argue that it is crucial to understand the inherent anonymity implications of stream-based communication itself, independently of other system properties such as latency or cover traffic.
-This is because stream-based communication is at the core of modern online communication.
-Our work therefore aims to shed light on basic, inherent anonymity properties of stream-based communication that are relevant not only for onion routing systems but also to other ACNs aiming to support stream-based communication.
-One step towards understanding such relations is to measure or estimate the achieved anonymity.
-While a plethora of anonymity metrics exists for many different applications [DBLP:journals/csur/WagnerE18], these are not primarily useful for understanding the (asymptotic) implications of stream-based communication.
-Our goal therefore is to quantify the anonymity level of streams in a way that helps understand inherent factors which increase or reduce anonymity.
+2 Related Work
+The study of anonymous communication began in 1981, when Chaum introduced the concept of a mix [chaum1981untraceable].
+The key idea of achieving unlinkability through independent messages of uniform size, making them indistinguishable to observers, has since become foundational for mixnet designs, including modern systems like Loopix [piotrowska2017loopix] and Nym [diaz2021nym].
+The anonymization of streams of data gained prominence with the introduction of onion routing [syverson1997anonymous], which became a core design principle in Tor [dingledine04tor].
+Other low-latency anonymity networks have since adopted this approach [chen2015hornet, DBLP:journals/popets/ChenP17].
+[…]
+3 Progressive Pruning
+In this section, we introduce progressive pruning, a novel methodology designed to analyze the anonymity implications of data streams in anonymous communication networks (ACNs).
 3.1 Adversary Model
-Discussing and quantifying anonymity only makes sense with regard to a concrete adversary model.
 In order to evaluate the anonymity properties of stream-based communication, we assume that an attacker can monitor both ends of the streams.
 That is, the attacker can observe when packets of each stream are sent and received.
-This setting therefore focuses on traffic analysis attacks in which the attacker compares temporal traffic patterns at both ends of the streams, trying to identify pairs of communication.
-In the context of stream-based communication the primary attack is an intersection attack [berthold2001disadvantages], which will therefore be the main concern of this work.
+This setting, therefore, focuses on traffic analysis attacks in which the attacker compares temporal traffic patterns at both ends of the streams, trying to identify pairs of communication (see Figure 1).
+In the context of stream-based communication, the primary attack is an intersection attack [berthold2001disadvantages], which will therefore be the main concern of this work.
 The idea behind such attacks is to exploit the knowledge of multiple messages belonging together by intersecting their anonymity sets in order to identify the communication partners.
 Our adversary model is inspired by a global passive adversary, which appears to be a sensible assumption today [nsa-buys-data], and deliberately abstracts from any concrete underlying communication system to capture the anonymity properties inherent to streams.
-2.3 Related Work
-The long research history of anonymous communication begins in 1981, when Chaum introduced the concept of a mix [chaum1981untraceable].
-Even in this early work, the central idea to achieve unlinkability revolves around the notion of independent messages that are uniformly sized and thus indistinguishable for an outside observer.
-This assumption has since become a core principle of mixnet designs, up to modern mixnet systems like Loopix [piotrowska2017loopix] and Nym [diaz2021nym].
-The idea to anonymize streams of data was first followed in the context of ISDN telephony [pfitzmann1991isdn].
-The system primarily exploited the fixed-rate nature of the ISDN network to achieve traffic analysis resistance.
-With the introduction of onion routing [syverson1997anonymous], the stream notion gained further attention, resulting in its adoption as a primary design principle in Tor [dingledine04tor].
-Other low-latency anonymity networks have since continued to utilize this approach [chen2015hornet, DBLP:journals/popets/ChenP17].
-However, streams of data being transferred in a low-latency fashion bring with them increased susceptibility to traffic analysis attacks, in particular intersection attacks [berthold2001disadvantages].
-This is also the fundamental approach followed by our progressive pruning methodology, whereas our work aims specifically at making the impacts of specific stream properties comprehensible.
-A version of this attack that instead operates on the basis of probabilities has been put forward as the statistical disclosure attack.
-In general, the field of traffic analysis attacks has been the subject of many research studies.
-Correlation techniques that have been proposed range from simple approaches like packet counting [DBLP:conf/ih/BackMS01] through information-theoretic estimates like cross-correlation coefficients [levine2004timing] and mutual information [DBLP:conf/pet/ZhuFGBZ04] to advanced matching techniques based on deep learning [DBLP:conf/ccs/NasrBH18].
-While these methods constitute relevant real-world attacks, they are not necessarily useful to understand the essence of anonymity within stream-based communication, as influenced e.g.
-by the length of streams or the number of simultaneous streams.
-We therefore propose an alternate methodology that supports establishing an understanding of how different inherent stream properties influence anonymity.
-[…]
-3 Progressive Pruning
+Figure 1: For some user (yellow), an attacker aims to identify the communication partner (red or blue) using traffic analysis.
 3.2 Overview
-We now introduce progressive pruning, a new methodology to analyze the anonymity implications of streams of data, as they are transferred via an ACN.
-We understand a stream as a reliable, unidirectional sequence of messages that are exchanged between the same pair of sender and receiver, over the same communication channel (e.g.
-through an ACN).
-Streams do not necessarily need to be order-preserving, however, messages need to be delivered within some reasonably bounded time.
-This transfer of streams of data comes with an increased vulnerability against intersection attacks.
-In such attacks, an attacker exploits the fact that multiple messages belong to the same communication partners by intersecting the anonymity set to reduce the number of candidates, keeping only candidates that are in line with all observations.
-Since the vulnerability to intersection attacks is what fundamentally differentiates streams from single messages, we design our method similarly.
-This way, we can identify anonymity effects that primarily stem from the stream characteristics, compared to the exchange of a single message.
-The high-level goal of progressive pruning is as follows: Given a network trace of stream messages along with observation timestamps at its source and destination, we want to calculate an anonymity estimate for each of the contained streams.
-For this, progressive pruning builds upon the well-known notion of anonymity sets.
-However, it adds a time component: Initially, there may be a broad set of candidates, so the first message may have a large anonymity set.
-However, as communication proceeds over more messages, the anonymity set shrinks as candidates drop out.
-This may be the case because at some point their observed behavior may not be compatible with other observations.
-The fact that the set of candidates is pruned progressively as the message streams advance is what gives our methodology its name.
-It should be noted that, as a consequence of the design, our anonymity metric does not provide an absolute value of system security or anonymity, which would be very hard or likely impossible.
+Progressive pruning builds on the well-known notion of anonymity sets—but evolves over time: Given a network trace of messages with observation timestamps at their source and destination, we model the core concept of an intersection attack and quantify anonymity sets per stream.
+Initially, a message may have a large anonymity set as there may be a broad set of candidates.
+However, as communication progresses and more messages are exchanged, the anonymity set shrinks as candidates are excluded.
+For example, the observed timing of events may conflict with other previously gathered observations, rendering them incompatible.
+Pruning the set of candidates progressively gives our methodology its name.
+We illustrate progressive pruning for a single message in Figure 2.
+We note that, as a consequence of the design, our anonymity metric does not provide an absolute value of system security or anonymity, which would be very hard or likely impossible.
 However, it can be used to analyze the relative changes to anonymity induced by different system parameters.
-3.2 Detailed Description
+3.3 Modeling Intersection Attacks
 Stream Model:
@@
 We further assume that there is no message loss, and the delay of each message (the time between being sent and received) is bounded between $[D_{min},D_{max}]$.
-Depending on whether we want to evaluate the anonymity level of senders or receivers, either (1) senders send all their messages to the same receiver, or (2) receivers receive all their messages from the same sender.
+Depending on whether we want to evaluate the anonymity of senders or receivers, either (1) senders send all their messages to the same receiver, or (2) receivers receive all their messages from the same sender.
 The attacker’s goal is to link senders and receivers.
-For clarity, we stick to the sender perspective from now on.
-Consequently, the attacker wants to find the corresponding receiver of each sender.
-The opposite direction is fully analogous.
-As our deliberately simple attacker represents a concrete attack instance, the resulting anonymity values can be interpreted as upper bounds on the achieved anonymity.
-Our assumptions regarding the stream model are not meant to mimic any existing real-world system, but rather allow investigating the inherent properties of stream communication.
-Figure 2: Anonymity set calculation for individual messages of a sending stream.
-For each message from the sender (yellow), the set of potentially corresponding receiver messages is determined based on the lower and upper network delay bounds $D_{min}$ and $D_{max}$.
-Receivers of messages within this window (red, blue) may be communication partners of the sender.
-The anonymity set of the whole stream is then calculated by intersecting the individual message anonymity sets.
+For clarity and without loss of generality, we focus on the sender’s perspective.
+Consequently, the attacker aims to identify the corresponding receiver for each sender.
+As our deliberately simple attacker represents a concrete attack instance, the resulting anonymity can be interpreted as upper bounds on the achieved anonymity.
+Our assumptions for the stream model allow investigating the inherent properties of stream communication but are not intended to replicate any real-world system.
 Terminology:
@@
 The relationship between sent and received messages can be characterized as follows: $\forall m_{S}\in M_{S}:\exists m_{R}\in M_{R}:m_{S}=m_{R}$.
-It is the adversary’s goal to find this mapping, i.e.
-determine the receiver of a sent message, breaking sender-receiver unlinkabilty.
+It is the adversary’s goal to find this mapping, i.e., determine the receiver of a sent message, breaking sender-receiver unlinkabilty.
 Furthermore, let $T_{S}[m]$ be the time when message $m$ was sent, and $T_{R}[m]$ the time when message $m$ was received.
@@
 We therefore define it as follows:
-$\Psi_{MM}[m_{S}]$ =$=\{m_{R}$ $\in M_{R}\mid T_{R}[m_{R}]\in[T_{S}[m_{S}]+D_{min}~{},~{}T_{S}[m_{S}]+D_{max}]\}$
-Based on this, it is straightforward to define the message-receiver anonymity set $\Psi_{MR}$, of a message, denoting the set of potential receiver candidates, a subset of $R$:
+$\Psi_{MM}[m_{S}]=\{m_{R}\in M_{R}\mid T_{R}[m_{R}]\in[T_{S}[m_{S}]+D_{min}~{},~{}T_{S}[m_{S}]+D_{max}]\}$
+We can use this to define the message-receiver anonymity set $\Psi_{MR}$ for a message, which represents the potential receiver candidates and forms a subset of $R$:
 $\Psi_{MR}[m_{S}]=\{R[m_{R}]~{}\forall~{}m_{R}\in\Psi_{MM}[m_{S}]\}$
+Figure 2: Progressive Pruning: For each sender message (yellow), potential receiver messages are identified based on the network delay bounds $D_{min}$ and $D_{max}$.
+Receivers of messages within this window (red, blue) are considered as candidates.
+The stream’s anonymity set is determined by intersecting these message anonymity sets over time.
 Anonymity set of multiple messages:
-We now extend these anonymity set definitions to multiple messages from the same sender (a stream).
+We now extend these anonymity set definitions to multiple messages from the same sender—a stream.
 By doing so, we want to capture the anonymity set of senders as a whole, taking all their messages into account.
@@
 For example, if a sender sent 10 messages in a short period of time, but a receiver received only one message that falls into the message-message anonymity sets of all sent messages, the simple intersection approach would not exclude that receiver although it obviously cannot be a candidate because it received too few messages.
-We solve this issue by formulating the following assignment problem: Let $s$ be a sender of messages $M_{S}[s]$ and $r$ a receiver of messages $M_{R}[r]$.
-Then, $r$ can only be a candidate as the receiver of the messages from $s$, if there exists a mapping $\sigma:M_{S}[s]\to M_{R}[r]$ that assigns each sent message to a corresponding received one and meets the following two requirements: (1) If $\sigma(m_{S})=m_{R}$, then $m_{R}\in\Psi_{MM}[m_{S}]$ (the messages potentially correspond with respect to the delay constraints), and (2) $\sigma$ is injective (each received message is assigned at most once).
-Given the sets of sent and received messages, it is possible to determine whether such a mapping exists, e.g.
-in a brute-force manner.
-Moreover, in Section 3.4 we will present an efficient algorithm to solve this task in linear time.
+We address this issue by formulating the following assignment problem: Let $s$ be a sender of messages $M_{S}[s]$ and $r$ a receiver of messages $M_{R}[r]$.
+Then, $r$ can be a candidate receiver for the messages from $s$ only if there exists a mapping $\sigma:M_{S}[s]\to M_{R}[r]$ that assigns each sent message to a corresponding received one and meets the following two requirements: (1) If $\sigma(m_{S})=m_{R}$, then $m_{R}\in\Psi_{MM}[m_{S}]$ (messages potentially correspond w.r.t.
+the delay constraints), and (2) $\sigma$ is injective (each received message is assigned at most once).
+Given the sets of sent and received messages, it is possible to determine whether such a mapping exists, e.g., in a brute-force manner.
+In Section 3.5, we will present an efficient algorithm to solve this task in linear time.
 If a mapping that meets the the aforementioned requirements can be found, we say that $r$ is a candidate for $s$.
@@
 Additional messages in the stream, however, add additional constraints that can be exploited by the adversary.
-As such, more messages will likely result in a shrinking anonymity set (if the streams of communication are not synchronized).
-At the same time, the number of observed messages that are necessary to break anonymity can be regarded as an indicator for the costs an adversary has to invest for the attack.
+As such, more messages will eventually decrease the anonymity set.
+The number of observed messages that are necessary to break anonymity can be regarded as an indicator for the costs an adversary has to invest for the attack.
 Another observation that can be made is that the number of messages sent by a sender is highly relevant for the resulting anonymity set.
-In particular, if the streams differ in their number of messages, then the stream length is useful for an adversary to differentiate the senders.
-While this comes at no surprise, it has an interesting implication: A sender that sends more data than anyone else does not retain any anonymity at all.
+In particular, if the streams differ in their number of messages, then the stream length is useful for an adversary to differentiate senders.
+While this comes at no surprise, it has an interesting implication: A sender that sends more data than anyone else does not retain any anonymity.
 In general, the shorter the communication is in comparison to the others, the more likely it is that a sender can “hide” among the other senders.
-We formally verified this important relationship and present the analysis in the appendix.
+We also verified this important relationship formally [progressive-pruning-arxiv].
 Additionally, the absolute latency value of messages does not directly make a difference.
 Instead, an attacker only benefits from a low deviation of latency (jitter) because it makes the traffic more predictable, allowing use of a smaller window.
-3.4 Algorithm: Matching Candidate Messages
+3.5 Algorithm for Matching Candidate Messages
 Here we present an efficient algorithm that can be used to determine whether a receiver $r\in R$ is a candidate for sender $s\in S$, based on their sets of sent and received messages, as defined in Section 3.3.
@@
 We do so by comparing $\psi(m_{S},r)$ to the (previous) message anonymity set $\psi(m_{S}^{\prime},r)$ of the previous message $m_{S}^{\prime}$.
-$m_{S}$ can mapped if and only if there is at least one valid candidate message.
+$m_{S}$ can be mapped if and only if there is at least one valid candidate message.
 The number of candidates (cand) is determined as the sum of leftover unused candidate messages from the previous anonymity set $\psi(m_{S}^{\prime},r)$ that are also present in $\psi(m_{S},r)$, which we call overlap, plus the number of new messages $|\psi(m_{S},r)\setminus\psi(m_{S}^{\prime},r)|$, which we call new.
@@
 Otherwise, this is not the case.
-3.5 Implementation: ppcalc
-We implement progressive pruning and make it available as an open-source project called ppcalc  [^1] https://github.com/cdoepmann/ppcalc, implemented in the Rust language.
+3.6 Implementation
+We implement progressive pruning and make it available as an open-source project called ppcalc  [^1] https://github.com/cdoepmann/ppcalc, implemented in Rust.
 ppcalc consumes network traces consisting of messages between senders and receivers, and their timestamps.
@@
 It supports the calculation of sender anonymity as well as receiver anonymity.
-During the implementation of ppcalc, great care was taken to achieve high scalability that enables ppcalc to process communication traces not only from relatively small, synthetic scenarios as considered in this section, but also traces from real-world networks like Tor, at scale.
+During the implementation of ppcalc, great care was taken to achieve high scalability that enables ppcalc to process communication traces not only from relatively small, synthetic scenarios but also traces from real-world networks like Tor, at scale.
 For the experiments we will later present in Section 5, ppcalc successfully managed to process network traces of approximately 720 GB within 74 hours on a 12 CPU cores machine.
-4 Understanding Anonymity Over Time
-Progressive pruning, as presented in Section 3.2, explicitly focuses on the evaluation of anonymity within stream-based communication.
-That is, it takes into account how the anonymity set changes over the course of a data transfer that consists of a sequence of messages.
-In order to understand the implications of this measurement approach, we therefore analyze how anonymity changes with an ongoing stream of messages.
-This can also be understood as a development over time as the single messages are generally sent sequentially.
-Note that progressive pruning does not require any specific temporal characteristics but works the same if all messages in a stream are sent simultaneously.
-For this first analysis, we rely purely on stochastical methods in order to learn about the time behavior that is inherent to our model.
-In order to do so, we first consider the following very simple model of communication over an anonymous communication network: For now, we assume that all users send their messages at the same rate.
-This assumption allows us to discretize the system into synchronized rounds, assuming that one round corresponds to the transmission of exactly one message by each user that is active during that round.
-Although this is a simplification, we argue that is still close to the overall model behavior due to its window semantics.
-Accordingly, the window size that defines which messages belong to the anonymity set corresponds to one round.
-Making use of the terminology from Section 3.2, messages that are sent in the same round form a common message-message anonymity set.
-More intuitively, senders belong to the same anonymity set if they send data at the same time (within the same window/round).
-This matches our intuitive understanding of traffic analysis, as users sending messages at clearly different times are easy to tell apart.
-For simplicity, we further assume that all senders transfer the same number $L$ of messages, in $L$ consecutive rounds.
-Let us consider a single sender $s$ within a busy anonymity system, starting to send its messages at round 00.
-Then, the transfer will last $L$ rounds.
-Moreover, let $N$ be the overall number of (independent) senders active in the system at this first round.
-For each round $t$, we want to find the anonymity set size $|\Psi_{SR}[s]|$, which we abbreviate as $A_{t}$ here.
-Obviously, $A_{0}=N$.
-Since our sender $s$ starts sending while the system is already in use by other senders, those senders will finish sending their messages at different times.
-Let $E_{i}$ be the time when sender $i$’s transfer ends (more precisely, the first round in which sender $i$ does not send a message anymore).
-Then, $E_{s}=L$ because $s$ sends for $L$ rounds and starts at round 00.
-In this scenario, the anonymity set will only shrink when other senders stop sending at some time.
-Note that users starting to send later do not have an impact on the anonymity set because they have not been candidates before and thus cannot be valid anonymity set members later.
-The anonymity set cannot grow over time because it corresponds to the intersection of observations over time.
-The anonymity set size is therefore determined by the number of senders that have stopped sending.
-Let $\bar{A}_{t}$ be the number of senders that have already finished their transfer before round $t$.
-Then the anonymity set size is $A_{t}=N-\bar{A}_{t}$, and $\bar{A}_{0}=0$.
-Without loss of generality, let us assume that $E_{s}$ corresponds to $E_{0}$, simplifying the following formulations.
-Since all senders transfer the same number $L$ of messages, and are independent from each other, we can calculate the probabilities for certain end times as follows:
-$P(E_{i}=1)=\dots=P(E_{i}=L)=\frac{1}{L}~{}~{}~{}\text{for $0<i\leq N$}$
-Consequently, we can accumulate the probabilities as follows (note that the events $\{E_{i}=t\}$ are all disjoint):
-$P(E_{i}\leq 0)=0\qquad P(E_{i}\leq L)=1\qquad P(E_{i}\leq t)=t\cdot\frac{1}{L}$
-The relationship between $\bar{A}_{t}$ and the different $E_{i}$ is as follows: For a fixed round $t$, $\bar{A}_{t}=t$ iff.
-$E_{i}\leq t$ for $x$ different $i$, and $E_{i}>t$ for the remaining $(N-x)$ $E_{i}$.
-Due to the fact that the events $\{E_{i}\leq t\}$ are all independent of each other, we can model each of them as a Bernoulli experiment with success probability $p=t\cdot\frac{1}{L}$.
-For each round $t$, $\bar{A}_{t}$ is therefore binomially distributed with $n=N$ and $p=t\cdot\frac{1}{L}$:
-$P(\bar{A}_{t}=x)=\binom{n}{x}\cdot p^{x}\cdot(1-p)^{n-x}$
-Consequently, the anonymity set size is distributed as follows:
-$P(A_{t}=x)=P(\bar{A}_{t}=N-x)=\binom{N}{N-x}\cdot p^{N-x}\cdot(1-p)^{x}\qquad\square$
-We can draw interesting insights about the anonymity set size as modeled by our progressive pruning. In particular, if we calculate the expected value for the anonymity set size in round $t$, we get:
-$E[A_{t}]=E[N-\bar{A}_{t}]=N-E[\bar{A}_{t}]=N-n\cdot p=N-N\cdot\left(t\cdot\frac{1}{L}\right)$
-As a consequence, we can see that the expected anonymity set size decreases linearly/anti-proportionally over time.
-Put differently, the longer the transfer lasts, the weaker the achieved anonymity gets.
-This observation clearly underlines the significance of the data stream length: The shorter the communication is, the more likely it is to hide among other senders.
-Also note that the simple system setup described here (same-length senders with independent start times) is equivalent in principle to another relevant scenario to consider in which all senders start sending at the same time but have independently selected transfer lengths.
 4 Anonymity Factors for Streams
-{subcaptionblock}
-.3 Refer to caption {subcaptionblock}.3 Refer to caption {subcaptionblock}.3 Refer to caption
-Figure 3:
-Figure 4:
-Figure 5:
-Figure 6: Influence of different stream properties on stream anonymity set size.
+[…]
+[…]
+[…]
+Figure 3: Influence of different stream properties on the anonymity set size.
 We now want to gain initial insight on the general factors that influence anonymity in the case of stream communication.
@@
 Note that this investigation complements existing work like [wright2003defending] and [johnson2013usersrouted] in that these focus on anonymity over time due to user behavior, while we consider anonymity properties of the streams themselves.
-Our general approach is as follows: We simulate communication between abstract senders and receivers with different scenario parameters.
-This simulation of communication streams is carried out on a high abstraction level—the communication properties are defined only by general parameters like stream length, latency, and number of senders and receivers.
+To this end, we simulate communication between senders and receivers with different scenario parameters such as stream length, latency, and number of senders and receivers.
 We then apply progressive pruning and use the calculated anonymity set sizes as estimates for anonymity.
@@
 For each of the parameters, we carry out a series of simulations in which we only vary this parameter.
-Our experiments reveal and verify the following relationships: Firstly, the length of a stream, in relation to other streams in the network, is of crucial importance for the achieved anonymity (cf.
-Figure 6).
-While short streams can achieve large anonymity sets, longer streams serve as cover traffic for the whole network but lose their own anonymity.
+In the following, we describe our results, revealing and verifying several relationships.
+Stream Length
+Firstly, the length of a stream, in relation to other streams in the network, is of crucial importance for the achieved anonymity (Figure 3).
+The results reveal that anonymity sets increase with growing stream length variation.
+At first sight, this is counterintuitive because more heterogeneous stream lengths may give an attacker a way to distinguish between streams.
+However, what can be observed here is the effect of cover traffic: Longer streams act as cover traffic for shorter streams, enhancing their anonymity while sacrificing their own.
+Due to the fact that a single long stream can act as cover traffic for several shorter streams, the average anonymity is increased.
 This underlines the effectiveness of cover traffic, either implicitly through a large user base, or explicitly using designated cover traffic.
-Secondly, a low number of receivers compared to the number of senders is beneficial for anonymity because it helps streams to blend in with each other (Figure 6).
-And thirdly, homogeneity of streams is also beneficial for anonymity.
-(Figure 6).
-While these results come as no surprise, they express the soundness and utility of predictive pruning to analyze anonymity properties of streams, also verifying these relationships based on a well-defined stream model.
+Network Topology
+We analyze the impact of network topology by varying the ratio between senders and receivers.
+With fewer receivers, each receiver handles the streams of more senders, leading to a higher-density network.
+We find that a low number of receivers compared to the number of senders is beneficial for anonymity because it helps streams to blend in with each other (Figure 3).
+This is in line with our expectation because a larger number of receivers leads to an increasingly sparse network, in which there are fewer candidates an attacker has to differentiate.
+However, a too low number of receivers is problematic for our notion of anonymity as well because the total number of receivers constitutes an inherent upper bound for the anonymity set size.
+Network Latency
+Lastly, we consider the case of varying network lateny (jitter).
+The results in Figure 3 indicate a small increase in anonymity with growing network jitter.
+There are two main reasons for this: Firstly, more network jitter generally makes the network less predictable for an attacker.
+And secondly, he introduction of jitter for the single messages effectively results in the streams spanning an increased period of time during transmission.
+If streams are active for a longer time, they will be active with more other streams simultaneously, leading to an increased anonymity set.
+All in all, these results come as no surprise, but express the soundness and utility of predictive pruning to analyze anonymity properties of streams, also verifying these relationships based on a well-defined stream model.
 5 Case Study: Comparing Tor Instances
-In the previous section, we demonstrated the behavior of progressive pruning for the analysis of stream-based communication, and inferred general factors that are relevant for the achieved anonymity.
-The considered scenarios, however, were synthetic examples suitable primarily for investigating and verifying our metric as well as general, abstract trends.
-In contrast, we now go further and aim to apply our analysis approach to the most widely used anonymity network as of today—the Tor network.
-By doing so, we want to give a practical example of how progressive pruning can be used to research the anonymity properties of large-scale anonymity systems providing stream communication.
-As laid out in Section 3, progressive pruning is not suitable for quantifying anonymity as an absolute value, but rather as a relative measure for comparing anonymity.
-We therefore refrain from quantifying “the anonymity level of Tor”.
-The quantification can, however, be used to compare different instances of the network, hinting at deviations in achieved anonymity.
-We use this approach to investigate the implications of one of the future changes Tor will face, namely network growth.
-Over time, the Tor network has grown significantly and it is expected to continue this trend in the future [DBLP:conf/IEEEares/DopmannT23].
-While this development indicates the increasing significance and the growing user base of Tor, it does not come without challenges.
-Upcoming issues should be spotted early, so they can be taken into consideration when designing future network iterations.
-In [DBLP:conf/IEEEares/DopmannT23], the authors introduced a methodology to investigate upscaled instances of the Tor network, with regard to different topological properties, also finding initial evidence that purely horizontal growth of the Tor network can lead to suboptimal performance characteristics.
-We now aim to leverage our progressive pruning methodology to analyze the scaling impact on anonymity provided by the network.
+So far, we demonstrated how progressive pruning can be used for analyzing stream-based communication and identifying factors relevant to anonymity.
+The considered scenarios, however, were primarily synthetic examples to verify our methodology and observe trends.
+Now, we apply this approach to the widely used Tor network, providing a practical example of how progressive pruning can assess the anonymity properties of large-scale systems.
+As discussed, progressive pruning quantifies anonymity relatively.
+Instead of defining “Tor’s anonymity level,” we compare network instances to identify deviations in anonymity, focusing on the implications of network growth.
+The Tor network has grown significantly over time and it is expected to continue this trend.
+This growth, however, presents challenges for future network designs.
+Previous work [DBLP:conf/IEEEares/DopmannT23] analyzed scaled Tor instances, revealing that purely horizontal growth may harm performance.
+Here, we leverage progressive pruning to explore the impact of scaling on network anonymity, which is necessary as adversaries become stronger.
 5.1 Methodology
-Our general approach is to estimate the achieved anonymity based on simulations of the network.
-For this, we first simulate the Tor network in order to obtain a set of circuits that were created, along with traces of the exchanged messages (cells) and their timestamps.
-We then compute the anonymity level according to progressive pruning.
-This process is first carried out for the original (unscaled) Tor network.
-Using torsynth [DBLP:conf/IEEEares/DopmannT23], we then generate two synthetic but representative Tor networks that were scaled either vertically or horizontally by +100%.
-We also evaluate these networks using the aforementioned workflow and then compare the achieved anonymity values.
-In order to avoid sampling artifacts, we refrain from down-sampling the networks for simulation.
-Instead, we carry out all experiments at scale, which constitutes a nontrivial engineering challenge to scalability of the involved software components.
+Our general approach is to quantify anonymity sets based on message traces obtained from network simulations.
+Specifically, we simulate the Tor network to obtain a set of circuits, along with traces of exchanged messages (cells) and their timestamps.
+We then compute anonymity sets over time using progressive pruning.
+This process is first applied to the original (unscaled) Tor network.
+In addition, we generate two synthetic but representative Tor networks scaled either vertically or horizontally by +100% using torsynth [DBLP:conf/IEEEares/DopmannT23].
+We conduct network simulations on these scaled networks and apply progressive pruning to the resulting traces to quantify anonymity sets as well.
+In order to avoid sampling artifacts, all experiments are conducted at scale, which presents a nontrivial engineering challenge for the scalability of the involved software components.
 We base our experiments on the network consensus from 2023-10-01 at 00:00, and simulate one hour of full network activity.
 Tor Flow Simulation:
-Given a specific snapshot of the Tor network (consensus), the network simulation step needs to generate a trace of network activity within the Tor network.
-In particular, this involves generating circuits and using them to transport data cells as would be the case in the real network.
-There are different approaches towards simulating the Tor network, trading between computational effort and realism.
-The most realistic simulations can be carried out using the shadow simulator which is able to run the unmodified Tor client.
-However, running such simulations at scale 1:1 (or even 2:1 in the case of up-scaled networks) was infeasible.
-While shadow enables a very fine-grained simulation to investigate performance and other networking effects, we are particularly interested in how circuits and cells are distributed across the network because these are the factors for anonymity as considered by our metric.
-We therefore choose a more abstract simulation approach, similar to the Tor path simulator (TorPS [johnson2013usersrouted]).
-TorPS simulates circuit generation by Tor users over time and has been used before to evaluate how anonymity changes when users construct many consecutive circuits.
+Given a specific snapshot of the Tor network (consensus), the first step is to generate a trace of Tor network activity.
+In particular, this involves generating circuits and using them to transport data cells, simulating network communication.
+Simulating the Tor network involves a trade-off between computational effort and realism.
+The most realistic approach uses the shadow simulator [jansen12shadow], capable of running the unmodified Tor software.
+However, running such simulations at scale 1:1 is infeasible.
+While shadow enables fine-grained simulation to investigate performance and other networking effects, we focus on the distribution of circuits and cells across the network, because these factors impact anonymity as considered by progressive pruning.
+We therefore adopt a more abstract simulation approach, similar to the Tor path simulator, TorPS [johnson2013usersrouted].
+TorPS simulates circuit generation over time and has been used to study anonymity changes before.
 However, TorPS suffers from severe scalability issues.
 Furthermore, it does not simulate user behavior at the packet level and thus cannot be used to generate the network traces we require.
-We therefore propose a novel simulator software which we call Tor Flow Simulator (TorFS) and make available as an open source project [^4] https://github.com/cdoepmann/torfs, TorFS inherits much of TorPS’s functionality.
-However, it is implemented in the Rust programming language with a focus on scalability, and adds the missing pieces needed to simulate Tor user behavior at the packet level.
-Most importantly, it applies the modeled circuit selection rules as well as the users’ circuit management.
-It also supersedes TorPS’s simplistic user model, replacing it with the state-of-the-art model introduced in [DBLP:conf/ccs/JansenTH18] and used by other publications in the field since then [DBLP:journals/popets/PahlAKE23, DBLP:conf/ccs/RochetWJMP20, DBLP:conf/uss/JansenVS19].
-As in [DBLP:conf/ccs/JansenTH18], user behavior is now modeled as a probabilistic process (an exponential distribution) with parameters derived from real-world measurements on the Tor network.
+We developed the Tor Flow Simulator (TorFS), an open-source simulator [^4] https://github.com/cdoepmann/torfs designed to simulate Tor user behavior at the packet level.
+TorFS inherits much of TorPS’s functionality.
+However, it is implemented in Rust with a focus on scalability and adds the missing pieces, including the modeled circuit selection rules and users’ circuit management.
+TorFS also supersedes TorPS’s simplistic user model, replacing it with the state-of-the-art model [DBLP:conf/ccs/JansenTH18].
+That is, user behavior is now modeled as a probabilistic process (an exponential distribution) with parameters derived from real-world measurements on the Tor network.
 Moreover, while TorPS simulates only the generation and management of circuits, TorFS also simulates the circuits themselves to generate network traces of the individual messages (cells).
 For this, it again makes use of the probabilistic model from [DBLP:conf/ccs/JansenTH18].
-Consequently, TorFS uses the stream model as well as the packet model from [DBLP:conf/ccs/JansenTH18] which are Markov chains to mimic real traffic behavior as observed on the real Tor network.
-At the same time, TorFS does not constitute a full-fledged network simulator like shadow or ns-3.
+Consequently, TorFS uses the stream model as well as the packet model from [DBLP:conf/ccs/JansenTH18], which mimic real traffic behavior as observed on the real Tor network.
+As a result, TorFS outputs message traces ready for analysis with ppcalc.
+Please note, however, that TorFS does not constitute a full-fledged network simulator like shadow or ns-3.
 It abstracts from networking effects in that cells entering the network are output after a constant latency at their destination.
-This makes it feasible to simulate the network at scale.
-Moreover, we argue that this is in line with our approach of focusing on the inherent anonymity effects within the Tor network as influenced by its topology.
-TorFS outputs message traces ready for analysis with ppcalc.
-We only focus on traffic that is directed from the Tor exit to the client because this constitutes the majority of Tor traffic.
-We apply the following mapping from simulated Tor behavior to the abstract message model from Section 3: Each cell that occurs in the simulation corresponds to a message in the abstract model.
-Multiple cells that form a data stream in the simulation also constitute a single stream of messages in the sense of our abstract model.
-The abstract sender of these messages is the according simulated exit.
-For each stream, we define an abstract receiver as the entity that receives all of this stream’s messages.
+Applying Progressive Pruning to Tor:
+In our evaluation, we focus on traffic that is directed from the exit relay to the client as this constitutes the majority of Tor traffic.
+We apply the following mapping from simulated Tor behavior to the message model from Section 3: Each cell that occurs in the simulation corresponds to a message in the model.
+Multiple cells that form a data stream in the simulation also constitute a single stream of messages in the sense of our model.
+The sender of these messages is the according simulated exit.
+For each stream, we define a receiver as the entity that receives all of this stream’s messages.
 This way, the stream-based anonymity can be evaluated independently for each stream in the network.
-Anonymity Computation:
-After generating message traces of the Tor network using TorFS, we analyze these using progressive pruning in order to estimate the anonymity level.
-For this task, we again used ppcalc (cf.
-Section 3.5).
-As laid out before, the anonymity estimate can be calculated in either one of two possible “directions”.
+After generating message traces of Tor, we analyze these using progressive pruning in order to calculate anonymity sets.
+For this task, we again use ppcalc.
+As laid out before, anonymity sets can be calculated in one of two possible “directions”.
 In our current experiment, we have a fixed (relatively small) set of exit relays that are sending data to many clients.
@@
 Stddev of anonymity set size 594.4 416.5 1,186.0
-{subcaptionblock}
-.45 Refer to caption {subcaptionblock}.45 Refer to caption
-Figure 7: Orig. (Refer to caption), vert. (Refer to caption), horiz. (Refer to caption)
-Figure 8: Original
-{subcaptionblock}
-.45 Refer to caption {subcaptionblock}.45 Refer to caption
-Figure 9: Vertically scaled
-Figure 10: Horizontally scaled
-Figure 11: (a) Overall anonymity distribution of streams per scenario, (b)-(d) Anonymity distribution of streams, grouped by stream length per scenario (from percentiles 0-10/black, to percentiles 90-100/orange).
-Our analysis is based on the streams’ anonymity set sizes and their distribution across the network.
-Table 1 displays fundamental statistics about the anonymity in each simulated network.
-As a first observation, we can see that, on average, the anonymity set sizes are relatively large.
-In the case of the unmodified Tor network, e.g., streams still remain with an average of more than 1,200 sender candidates (out of a total number of approximately 2,400 exits).
+Figure 4: Anonymity sets.
+Figure 5: Anonymity sets by stream length (percentiles).
+[…]
+[…]
+[…]
+Our analysis is based on anonymity set sizes and their distribution across the network.
+Table 1 displays basic statistics about the anonymity in each simulated network.
+As a first observation, we can see that the anonymity set sizes are on average relatively large.
+In the case of the original Tor network, e.g., streams still remain with an average of more than 1,200 sender candidates (out of approximately 2,400 exits).
 It is noteworthy that, with regard to the simple window-based attack progressive pruning mimics, anonymity of the vast majority of streams is not trivially broken.
-This observation is supported by the fact that—looking again at the original Tor network—out of 26 million streams, less than 1,000 are fully de-anonymized (resulting in an anonymity set size of 1).
-In order to better understand how anonymity is distributed within the network, we consider CDF plots of the achieved anonymity set sizes, displayed in Figure 11.
+This observation is supported by the fact that—looking again at the original Tor network—out of 26 million streams, less than 1,000 are fully de-anonymized, i.e., resulting in an anonymity set size of 1.
+In order to better understand how anonymity is distributed within the network, we consider CDF plots of the anonymity set sizes, displayed in Figure 5.
 Looking at the data from the original Tor network first, it becomes immediately apparent that anonymity is distributed very unevenly between the streams.
@@
 Having seen in Section 4 that a stream’s anonymity is heavily influenced by its length, we next investigate the relationship between stream lengths and anonymity in the simulated networks.
-Figure 11 contains CDF plots of the achieved anonymity set sizes again, but subdivides the streams into quantiles based on their length (number of cells).
-The resulting anonymity distributions clearly point out that stream length is again a decisive factor for the resulting anonymity.
+Figure 5 shows CDF plots of the anonymity sets, subdivided into quantiles based on their length (number of cells), from percentiles 0–10 (black), to 90–100 (orange).
+The resulting anonymity distributions clearly point out that stream length is again a decisive factor for anonymity.
 Long-lasting streams are massively disadvantaged and generally achieve only little anonymity.
@@
 These results again support our observations from Section 4: Long streams suffer from reduced anonymity but act as cover traffic providing stronger anonymity for shorter streams.
-Anonymity in Scaled Networks
-We leverage our simulation and analysis workflow to evaluate not only the Tor network as is today but also variations of the network that could result from future network growth of different characteristics.
-Our choice to evaluate scaled Tor networks is motivated by two goals: Firstly, it allows us to research the influence of network topology on the achieved anonymity.
-And secondly, understanding the trade-offs of different topological properties may provide important insights on how to shape and prepare for future network evolution.
+Anonymity in Scaled Networks:
+We leverage our simulation and analysis workflow to evaluate not only the Tor network as is today but also variations of the network that could result from future network growth with different characteristics.
+Evaluating scaled Tor networks allows us to investigate the influence of network topology on anonymity.
+Understanding the trade-offs of different topological properties may provide important insights on how to shape and prepare future network evolution.
 In addition to a snapshot of today’s Tor network, we therefore carried out simulations of the Tor network being scaled by +100% vertically as well as by +100% horizontally.
@@
 This is somewhat surprising because, essentially, the number of exits only directly defines the maximum anonymity set size which wouldn’t necessarily have to correspond linearly to the average anonymity set size.
-It is noteworthy that the standard deviation also doubles exactly.
+Even if the anonymity set sizes grow with horizontal scaling of the network, it has to be noted that the standard deviation grows massively, too.
+In our experiments, it also exactly doubles.
+We regard this as a problematic property because it means that anonymity is distributed much more unevenly among users of the network.
 Looking at the vertically upscaled network instead paints a rather different picture.
@@
 All in all, we can conclude from the simulations of the Tor network which we carried out that there are two fundamental factors for anonymity in the network: Firstly, an increased number of relays also increases the number of (topological) candidates for the anonymity set.
-And secondly, apart from the topological factor, the mere increase in network load also improves anonymity, albeit much less effectively.
+And secondly, the mere increase in network load also improves anonymity, albeit much less effectively.
 6 Conclusion
-In this work, we analyzed the inherent susceptibility of stream-based anonymous communication to traffic analysis attacks.
-In particular, we focused on intersection attacks because these constitute the class of attacks which streams are fundamentally vulnerable against, compared to independent messages.
-We introduced progressive pruning, a methodology to quantify the anonymity level of streams, with regard to the effectiveness of an intersection attack.
-With that, we contribute a new perspective: Instead of regarding the susceptibility to intersection attacks as a binary property, we enable a more differentiated view.
-The ability to compare the anonymity achieved by different system variants can be used for designing new ACNs, taking attackers’ cost into account.
-We carried out simulations to investigate the impact of different stream properties in tailored synthetic scenarios.
-Moreover, we demonstrated how progressive pruning can be applied to large systems like the Tor network, estimating anonymity factors with regard to network topology.
-Our results indicate that the susceptibility towards traffic analysis attacks of streams is primarily dependent on the length of streams, relative to each other.
-Longer streams suffer from worse anonymity because they can more easily be distinguished from others by considering their number of messages.
-At the same time, they serve as cover traffic, increasing anonymity for shorter streams.
-When looking at Tor as an example of stream-based anonymity networks, we observe different influences when scaling the network.
-On the one hand, scaling the network horizontally (adding relays) helps increase anonymity set sizes in the context of our anonymity estimate that takes the perspective of a global adversary who monitors incoming and outgoing traffic of the anonymity system.
+In this work, we analyzed the inherent vulnerability of stream-based anonymous communication to traffic analysis.
+We specifically focused on intersection attacks, which exploit fundamental weaknesses of streams compared to independent messages.
+To this end, we introduced progressive pruning, a methodology to quantify stream anonymity in the face of such attacks, enabling a differentiated assessment beyond a binary view of susceptibility.
+This approach facilitates comparing anonymity across system variants, aiding the design of new ACNs that account for attackers’ costs.
+We conducted simulations to examine the impact of stream properties in various scenarios.
+Our findings highlight that stream length relative to others significantly influences susceptibility: longer streams are more easily distinguished, reducing their anonymity, but also serve as cover traffic for shorter streams.
+Additionally, we demonstrated the application of progressive pruning to large real-world networks, By using Tor as a case study, we observed that horizontal scaling improves anonymity by increasing anonymity set sizes under a global adversary model, but anonymity is distributed very unevenly.
 On the other hand, when considering a vertically scaled network, we can see that an increased number of users is beneficial for anonymity.
-We argue that stream-based communication is of high practical relevance for online privacy and should therefore be a goal when designing future anonymity systems.
-While our work shows that it has inherent, negative consequences for anonymity due to being vulnerable to intersection attacks, our results may serve as a starting point in working around these limitations with the vision of combining stream-based communication with strong anonymity.
+Despite the inherent vulnerability to intersection attacks, our findings provide a methodology to systematically explore mitigation strategies increasing the difficulty faced by adversaries.
 […]